OpenStudy (anonymous):

3r/4=1/8 is r=1/6 right?

ganeshie8 (ganeshie8):

\(\large \color{Red}{\checkmark}\)

OpenStudy (luigi0210):

Yup, you are correct.

OpenStudy (luigi0210):

Pfft, gane and his fancy little check mark.

OpenStudy (anonymous):

Haa thank u

OpenStudy (anonymous):

2r/3=16 r= 24 right

OpenStudy (phi):

if you multiply both sides by 4/3 you get r = 1/8 * 4/3

OpenStudy (phi):

2r/3=16 multiply both sides by 3/2 r = 16*3/2

OpenStudy (phi):

\[ r = \frac{16\cdot 3}{2}= \frac{8\cancel{16}\cdot 3}{\cancel{2} 1} \]

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

how about -15= -3n/5 n would equal 25 right

OpenStudy (phi):

n= -5/3 * -15 - -5 * -5 = 25

OpenStudy (anonymous):

so i was right , right

OpenStudy (phi):

right

OpenStudy (anonymous):

thanks agin